OpenStudy (theflash12345):

How do I solve the inequality-2(w-7)+3>w-1

OpenStudy (welshfella):

first distribute the -2 over the parentheses

OpenStudy (welshfella):

can you do that?

OpenStudy (theflash12345):

Yes

OpenStudy (welshfella):

ok what do you get?

OpenStudy (tomboy213):

easy -2w+10>w-1

OpenStudy (tomboy213):

I combined like term 7+10 too

OpenStudy (welshfella):

@tomboy - why dont you let TheFlash try it himself?

OpenStudy (theflash12345):

I got -2w-11>w-1

OpenStudy (welshfella):

that is wrong

OpenStudy (welshfella):

-2(w - 7) + 3 = -2*w -2*-7 + 3 = -2w + 14 + 3 = -2w + 17

OpenStudy (welshfella):

so we have -2w + 17 > w - 1

OpenStudy (theflash12345):

I got that now

OpenStudy (welshfella):

can you continue from there Flash?

OpenStudy (theflash12345):

Yes thanks

OpenStudy (welshfella):

yw